OpenStudy (anonymous):

Simplify: The SQUARE ROOT OF: 25/28

OpenStudy (anonymous):

=5/2sqrt(7)

OpenStudy (saifoo.khan):

\[\Huge \frac{5}{2\sqrt{7}}\]

OpenStudy (anonymous):

The original problem was 5/7 * 35/28 all under the RADICAL

OpenStudy (saifoo.khan):

u did right.

OpenStudy (anonymous):

any difference in the answer?

OpenStudy (anonymous):

thanks~

OpenStudy (saifoo.khan):

Nope.